$SPAD/src/input dmp.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool dmp.output
)set message test on
)set message auto off
)clear all
-- Input generated from DistributedMultivariatePolynomialXmpPage
--S 1 of 8
(d1,d2,d3) : DMP([z,y,x],FRAC INT)
--R
--R                                                                   Type: Void
--E 1
--S 2 of 8
d1 := -4*z + 4*y**2*x + 16*x**2 + 1
--R
--R
--R                 2       2
--R   (2)  - 4z + 4y x + 16x  + 1
--R           Type: DistributedMultivariatePolynomial([z,y,x],Fraction(Integer))
--E 2
--S 3 of 8
d2 := 2*z*y**2 + 4*x + 1
--R
--R
--R            2
--R   (3)  2z y  + 4x + 1
--R           Type: DistributedMultivariatePolynomial([z,y,x],Fraction(Integer))
--E 3
--S 4 of 8
d3 := 2*z*x**2 - 2*y**2 - x
--R
--R
--R            2     2
--R   (4)  2z x  - 2y  - x
--R           Type: DistributedMultivariatePolynomial([z,y,x],Fraction(Integer))
--E 4

--S 5 of 8
groebner [d1,d2,d3]
--R
--R
--R   (5)
--R        1568  6   1264  5    6   4   182  3   2047  2    103      2857
--R   [z - ---- x  - ---- x  + --- x  + --- x  - ---- x  - ---- x - -----,
--R        2745       305      305      549       610      2745     10980
--R     2    112  6    84  5   1264  4    13  3    84  2   1772       2
--R    y  + ---- x  - --- x  - ---- x  - --- x  + --- x  + ---- x + ----,
--R         2745      305       305      549      305      2745     2745
--R     7   29  6   17  4   11  3    1  2   15     1
--R    x  + -- x  - -- x  - -- x  + -- x  + -- x + -]
--R          4      16       8      32      16     4
--R     Type: List(DistributedMultivariatePolynomial([z,y,x],Fraction(Integer)))
--E 5
--S 6 of 8
(n1,n2,n3) : HDMP([z,y,x],FRAC INT)
--R
--R                                                                   Type: Void
--E 6
--S 7 of 8
(n1,n2,n3) := (d1,d2,d3)
--R
--R
--R            2     2
--R   (7)  2z x  - 2y  - x
--RType: HomogeneousDistributedMultivariatePolynomial([z,y,x],Fraction(Integer))
--E 7
--S 8 of 8
groebner [n1,n2,n3]
--R
--R
--R   (8)
--R     4     3   3  2   1     1   4   29  3   1  2   7        9     1
--R   [y  + 2x  - - x  + - z - -, x  + -- x  - - y  - - z x - -- x - -,
--R               2      2     8        4      8      4       16     4
--R       2        1   2      2       1     2    2   1
--R    z y  + 2x + -, y x + 4x  - z + -, z x  - y  - - x,
--R                2                  4              2
--R     2     2     2   1     3
--R    z  - 4y  + 2x  - - z - - x]
--R                     4     2
--RType: List(HomogeneousDistributedMultivariatePolynomial([z,y,x],Fraction(Integer)))
--E 8
)spool

)lisp (bye) ———-

5

References^{[1] nothing}

6